OpenStudy (anonymous):

Without using (10), show directly that\[\left( 17.6^{-1} \right)^{8} = \left( 17.6^{-8} \right)\]

OpenStudy (anonymous):

what is (10)?

OpenStudy (anonymous):

The number 10 @RaphaelFilgueiras

OpenStudy (anonymous):

it's just the last part

OpenStudy (anonymous):

from the exponent rules of \(\left(a^m\right)^n=a^{m\times n}\) so if m=-1 and n=8, then m x n = -8... that's why \[\left( 17.6^{-1} \right)^{8} = \left( 17.6^{-8} \right)\]